jhonyy9 (jhonyy9):

- may exist pythagora's triplet formed by primes ? ,,pythagora's triplet same with 3-4-5 5^2 =3^2 +4^2 - or allways need contain,need being one even between ?

OpenStudy (mww):

Here's an interesting rule: for a Pythagorean triad a^2 + b^2 = c^2 where a < b < c you will find that b+c = a^2 if a is odd and 2(b+c) = a^2 if a is even. This can be a very interesting way to generate new triads. 3, 4, 5 --> 4+5 = 3^2 5, 12, 13 --> 12 + 13 = 5^2 7, 24, 25 --> 24 + 25 = 7^2 8, 15, 17 --> 2(15+ 17) = 8^2 9, 40, 41 --> 40 + 41 = 9^2 So the odd starting triads also have the property that c = b+1 Thus 2b + 1 = a^2. Now having established a (and thus a^2 is odd, this implies b must then be even), so your triad must have one even number. Similarly given c = a^2 - b = ODD - EVEN, you'll have two odd numbers and one even in your triad. If you start off with even triad, then 2(b+c) = a^2 and c = b+2 so 2(2b+2) = a^2 = 4(b+1) hence b must be odd and also c. So it holds all Pythagorean triads have two odd and one even component.

OpenStudy (mathmate):

@jhonyy9 unfortunately, there is always an even number >2, so cannot have a triad of all primes. Why always an even number? We know that odd^2=odd, even^2=even So if the triad is a^2+b^2=c^2, then odd^2+odd^2=odd+odd=even=even^2. and of course odd^2+even^2=odd+even=odd=odd^2. so there, no all prime triads!

OpenStudy (misty1212):

all triples are of the form \[m^2-n^2, 2mn, m^2+n^2\] primitive if \(gcf(m,n)=1\) and opposite parity so in fact not only is one even, it is divisible by 4

OpenStudy (misty1212):

in fact, if \(a,b,c\) is a primitive triple, one is divisible by 3, one is divisible by 4 and one is divisible by 5 not necessarily distinct ones

OpenStudy (mathmate):

@misty1212 Beautiful observation, and analysis. Thank you! so @jhonyy9 referring to previous post the closest as far as primes go, the best we can do is two primes out of three! "not necessarily distinct" means 3 and 4 can be divisors of the same member of the triad, as in 5, 12, 13.

OpenStudy (mathmate):

"not necessarily distinct" means, for example, 3 and 4 can be divisors of the same member of the triad, like 12 in (5, 12, 13), or 3 and 5 can be divisors of 15 in (8,15,17), while 4 is a divisor of 8.